OpenStudy (anonymous):

have to integrate (1/.2s+10)ds They get 5ln(.2s+10)-5ln10. Having limits of 0 and S. How did they get that?

OpenStudy (anonymous):

\[\int\limits_{0}^{s}(1/.2)s +10)ds\]?

OpenStudy (anonymous):

1/(.2s+10)ds

OpenStudy (anonymous):

(.2s+10)=u du=.2 ds ds=1/.2=5 \[5\int\limits_{0}^{s}1/u du\]

OpenStudy (anonymous):

-- correction ds=1/.2 du=5du